I need help calculating the finance charge and new balance using the previous balance method.
Previous balance = $179.32
Annual rate = 16%
Finance charge = $ ?
New purchases = $117.42
Payments/credits = $85.00
New balance = $ ?

I also need help calculating the unpaid balance, finance charge, and the new balance using the unpaid balance method.
Note: interest rate given as a monthly rate.
Previous balance = $179.32
Payments/credits = $85.00
Unpaid balance = $ ?
Monthly rate = 1.25%
Finance charge = $ ?
New purchases = $117.42
New balance = $ ?

Thank You. Once these frusterating questions are finished this completes my assinment.
Answer #1

Finance charge=179.32×(0.16÷12months)

=2.39

New balance= previous balance-payments/credit+finance change+New purchases

New balance=179.32−85+2.39+117.42

=214.13


answered by: mathpower
Answer #2

== PART 2 ==

Unpaid balance = 179.32-85=94.32

Finance charge = 94.32×(0.0125÷12)=0.09825

New balance = 179.32-85+0.09825+117.42=211.84

answered by: mathpower